Solve the following:
80 = 10x


Solve:
2/5x = -10

Found 3 solutions by dead_famous1991, Austin, mystic1306:
Answer by dead_famous1991(7)

1) 80 = 10x
=> 80/10 = x
=> 8 = x


2) 2/5x = -10
=> 2 = -50x
=> -1/25 =x

Answer by Austin(2)
Answer by mystic1306(20)

80=10x
First get the x by itself
to do this divide both sides by 10
80/10=8
10x/10= 1x or x
so..
8=x

2/5x= -10
get the x by itself by multiplying both sides by 5 to make the 5 a fraction put 5 over 1

2x/5 times 5/1 the fives cancel so you have 2x/1 or 2x
-10 times 5= -50
so
2x= -50
now divide both sides by 2
2x/2=1x or x
-50/2= -25
x=-25


just to show you:
2x/5 times 5/1= 2x
-10 times 5= -50
2x= -50
2x/2=x
-50/2= -25
x= -25